OpenStudy (doc.brown):

What's the determinant of a 1x1 matrix?

OpenStudy (doc.brown):

\[\left[ -4 \right]\]

OpenStudy (anonymous):

Ask Kai

OpenStudy (kainui):

Well I would just say that since: \[[-4]=-4*I_1\] so it's really just the scalar -4 multiplied by the identity matrix, and the determinant of an identity matrix is just 1. so... -4. You could think about it a little more in depth if you want too. For instance, when you break a nxn matrix determinant down into a bunch of (n-1)x(n-1) determinants.

OpenStudy (kainui):

Interesting question though, how did this come up? In your class or just curious? =)

OpenStudy (anonymous):

I like you in depth explanation haha

OpenStudy (kainui):

Well it's sort of vague and should let you think about it haha. I can always elaborate but I would rather have people ask questions about it based off what they think I mean. =P

OpenStudy (doc.brown):

I figured by cofactor expansion along the first row i would get +4 plus nothing else.

OpenStudy (doc.brown):

I was wondering if absolute values \(|-4|\) were just mini determinants.

OpenStudy (kainui):

Nope, definitely not. When I said that I was more or less thinking about turning a 2x2 matrix into a couple determinants of 1x1 matrices like this by expanding across the top row. \[\det \left[\begin{matrix}a & b\\ c & d\end{matrix}\right]=a*\det[d]-b*\det[c]\]

OpenStudy (doc.brown):

but then det[d] is just d. hmmm

OpenStudy (doc.brown):

Alright, you win.

OpenStudy (kainui):

Yup. =P